OpenStudy (leozap1):

Fan & Medal Is the relationship between the variables in the table a direct variation, an inverse variation, both, or neither? If it is a direct or inverse variation, write a function to model it. X 2, 5, 15, 20 Y 20, 15, 2, 2 A. Inverse variation; y = 10/x B. Neither C. Direct variation; y = x + 10 D. Direct variation; y = 10x

OpenStudy (anonymous):

@leozap1 please come back

OpenStudy (anonymous):

y has 2,2 the last two terms ?

OpenStudy (anonymous):

or 5,2 same with x line ?

OpenStudy (leozap1):

The last two terms for y is 2, 2 yes.

OpenStudy (anonymous):

ok so than what is your opinion what will be the righ choice ?

OpenStudy (anonymous):

right choice - sorry

OpenStudy (leozap1):

Direct variation; y = x + 10 ?

OpenStudy (anonymous):

do you can prove it too ?

OpenStudy (leozap1):

How?

OpenStudy (anonymous):

exactly this i said how you prove that y=x+10 is the right answer ?

OpenStudy (leozap1):

Wait, it only works for 5.

OpenStudy (anonymous):

X 2, 5, 15, 20 Y 20, 15, 2, 2 so than you see the first one case x=2 and y=20 so result y=10x second column 15=3x 3rd, 2 = ?15 4th 2=20/10 so what is the right answer than ?

OpenStudy (anonymous):

how for 5 work ?

OpenStudy (leozap1):

I would say y = x + 10, but that doesn't work for the last two.

OpenStudy (leozap1):

Because 5+10=15., but that's the only one that works on.

OpenStudy (anonymous):

how you say this 20 not equal 10+2

OpenStudy (anonymous):

for the first column

OpenStudy (leozap1):

2x10 = 20

OpenStudy (anonymous):

X 2, 5, 15, 20 Y 20, 15, 2, 2 1 column 20=2*10 2. 15=10+5 3. 2 = (15+5)/10 4. 2= 20/10 i think it in this way is right

OpenStudy (anonymous):

@leozap1 so i think there are every 4 choices one to one column try check it and will see

OpenStudy (anonymous):

so do you see it now ? first column x=2 y=20 is choice D. bc. 20=10*2 yes ?

OpenStudy (anonymous):

so second column what choice will be ?

OpenStudy (leozap1):

So the answer is B?

OpenStudy (anonymous):

for second column 15=5+10 why B. and not C. ?

OpenStudy (leozap1):

It doesn't work with every column, so it couldn't be C could it?

OpenStudy (anonymous):

please try understand me this choices are one to one for every column

OpenStudy (anonymous):

so for this second column is correct C. bc. 15=5+10

OpenStudy (leozap1):

X 2, 5, 15, 20 Y 20, 15, 2, 2 2 + 10 = 12 not 20 5 + 10 = 15 (it works) 15 + 10 = 25 not 2 20 + 10 = 30 not 2

OpenStudy (anonymous):

for 3rd column 2 and 15 will be choice B. and for the last one will be choice A.

OpenStudy (anonymous):

do you understand it now ?

OpenStudy (anonymous):

@ijlal your opinion about this please ?

OpenStudy (leozap1):

I can only choose 1 answer though...

OpenStudy (anonymous):

not possiblie bc there are every column true for one choice

OpenStudy (ijlal):

let me solve this i'll get back to you

OpenStudy (leozap1):

@ijlal What did you get?

OpenStudy (ijlal):

well i agree with @jhonyy9 every column gives a different variation

OpenStudy (leozap1):

Out of the given options for the answer, what do you think would be the correct one?

OpenStudy (leozap1):

A. Inverse variation; y = 10/x B. Neither C. Direct variation; y = x + 10 D. Direct variation; y = 10x

OpenStudy (ijlal):

you can only have 1 option as you're answer? it cant be that check again please ! i think you have to put choice respective to the columns like for the first column D is the answer

OpenStudy (leozap1):

Yea, I can only select one of the answers...

OpenStudy (ijlal):

i would say neither than option B would be your answer as every column has a different variation not 1 so B would be your answer :)

OpenStudy (leozap1):

That's what I started to think, it's kinda strange lol.

OpenStudy (ijlal):

yeah it was cause very column's variation was different that is why the answer ought to be neither :)